Restart Night #6 UT 2018-09-07

OSA: D.G. Huerta

Observers/ISA: B. Rothberg, J. Hill (Tucson)

Software: P. Kubanek

ISp: J. Riedl, G. Bechetti

Mtn Manager: J. Prothro

Details

UT 00:15 - Increased mount settling time for offset from 0 to 200 msec in iif.conf (although guider images weren't smeared, so this may not be the problem - I need to search through my old Chris Biddick emails as I think we've seen this before.).

UT 01:20 - Sky is mostly clear (much better than the past few nights), but one big thunderstorm building on the eastern boundary of Graham count may give us some trouble. Wind is light. T=10 degC with 70% humidity.

UT 01:55 - I've changed the sign of IDL LFBG agw_params[i].wfsc.x_pupil_shift_direction = 1.0 based on the dithered LUCI data from last night. And corresponding change for GCSL: WFS_pupil_shift_x -1.0 (And the sleep deprived person wonders why these two signs are apparently opposite ???)

UT 03:40 GCS collimating both sides in 1.4 arcsec seeing (with hardly any change from last night)

UT 03:45 restart GCS to use the latest patched version of today.

Apparent Vibration from Mirror Ventilation

Notice that "super noisy" vibration started at 2.875 UT -- just around the time when David turned on MV.

03.015 MV turned off, and vibration went back down, X Tilt rms is 122 and 35 mas.

03:16 Right MV only. X Tilt rms is 97 and 322 mas.

03:25 Left MV only. X Tilt rms is 434 and 52 mas.

03:50 Both MV back on. X Tilt rms is 449 and 318 mas.

Created IT 7394 - will add plots later.

Pupil Wobble measurements to align M3

using BS9183

Data from IDL in /home/lbto/data/20180907_lfbg and 20180907_rfbg

DX M3 at absolute Tip,Tilt -1200,400 arcsec (the old position).

03:55 pupil_collect, /OPEN, exp=10.0

The rotator Offset cancelled my preset, because it was the RotServicePosition command ?

After talking to Doug we learn that:
/OPEN means move the rotator at the engineering level when the telescope is not on an active preset. NOT /OPEN means use the normal Offset command to offset the rotator without cancelling the preset.

04:0x resend ACTIVE preset from IDL

press stopWFSing

04:11 pupil_collect, exp=10 without the /OPEN flag (running in two separate IDL windows for LFBG/RFBG)

seeing down to 0.7 on the right. SX collimation still around 1.2 arcsec.

04:22 Authorize MODS/MODS then back to LUCI/LUCI (trying to figure out why OSS isn't reading positions.conf) That did it.

04:27 DX M3 now at -865,-360. (Why we didn't find the DX star at the start of the night.) SX at -580,-160

04:30 ACQUIRE to recover pointing.

04:49 DX pupil_collect, exp=10.0 WITHOUT stopping WFS. This fails because IDL can't take its own wfs image when the WFS loop is running.

04:51 DX pupil_collect, exp=10.0 with wfs stopped.

Dreams for future improvements

  • pupil_collect needs a flag for reversing directions.
  • procedure needs an auto-repointing script to move M3 along with IE/CA in large steps (see sallanson email)

05:03 Moved SX M3 by +300 Tip (-280,-160 abs) but SX AdSec offline for a telemetry problem. David and Doug are working on the AdSec. Undid this change of M3.

05:04 Moved DX M3 by -300 Tip (-1165,-360 abs)

05:05 DX pupil_collect, exp=10.0 with wfs stopped. (This was the wrong way.)

05:18 Moved DX M3 by +500 rel Tip (-665,-360 abs)

05:18 DX pupil_collect, exp=10.0 with wfs stopped.

05:28 Moved DX M3 by +200 Tilt (-665,-160 abs)

05:28:58 DX pupil_collect, exp=10.0 with wfs stopped.

05:32 Moved SX M3 by +300 Tip (-280,-160 abs)

05:34 SX pupil_collect, exp=10.0 with wfs stopped, but it isn't collimating very well with IDL or GCS.

05:37 Moved DX M3 by +120 rel Tilt (-665,-40 abs)

05:37 DX pupil_collect, exp=10.0 with wfs stopped.

05:47 Moved DX M3 by +120 rel Tilt (-665,+80 abs) ADOPTED for RFBG

05:49 Moved SX M3 by -360 rel Tip (-640,-160 abs)

05:48:20 DX pupil_collect, exp=10.0 with wfs stopped.

05:5x SX pupil_collect, exp=10.0 with wfs stopped.

05:59 Moved DX M3 by +90 rel Tip (-575,+80 abs)

06:01 DX pupil_collect, exp=10.0 with wfs stopped.

06:03 Moved SX M3 by +120 rel Tip (-520,-160 abs)

06:04 SX pupil_collect, exp=10.0 with wfs stopped.

06:12 Moved SX M3 by -60 rel Tip (-580,-160 abs)

06:13 Moved SX M3 by +60 rel Tilt (-580,-100 abs)

06:14 SX pupil_collect, exp=10.0 with wfs stopped.

06:20 Moved DX M3 by -150 rel Tip (-725,+80 abs)

06:20 DX pupil_collect, exp=10.0 with wfs stopped.

06:25 Move SX M3 back to the original position (-580,-160) - satisfying in some sense that this is the best, but it still has ~7 pixels of wobble. Unclear why nearby values in Tip-Tilt space make a linear shift. Problem in pupil measurement? Some mechanical hysteresis?

06:26 SX pupil_collect, exp=10.0 with wfs stopped.

06:32 Moved DX M3 by +45 rel Tip (-680,+80 abs)

06:33 DX pupil_collect, exp=10.0 with wfs stopped.

(OSS restarted with the new DX M3 params the following afternoon.)

DX Pupil Wobble Stats (post facto analysis on 20200903
filename              Xavg     Yavg  Radius      pts
pupil_06:33:44.wobble 380.555 269.2925 12.4946058361 4
pupil_06:20:44.wobble 376.6175 266.57 16.126930117 4
pupil_06:00:22.wobble 378.3275 269.235 10.6472332087 4
pupil_05:48:20.wobble 379.0375 270.7775 8.61350973762 4
pupil_05:37:40.wobble 377.6525 270.595 14.0286697516 4
pupil_05:28:58.wobble 377.33 271.3725 11.9717949782 4
pupil_05:18:44.wobble 376.7425 269.7275 14.3525520379 4
pupil_05:05:31.wobble 376.0275 272.625 18.2306493302 4
pupil_04:51:38.wobble 375.415 271.765 16.5671723598 4

  • DX M3 Pupil Wobble Plot: (plot is of the latest 6 data sets)
    Screenshot at 2020-09-03 22-15-49.png

SX Pupil Wobble Stats (post facto analysis on 20200903
filename              Xavg     Yavg  Radius      pts
pupil_06:26:30.wobble 162.8525 215.7275 4.41723329699 4
pupil_06:14:37.wobble 162.9325 215.61 5.40740926877 4
pupil_06:04:19.wobble 162.9575 215.6675 4.34506041385 4
pupil_05:49:51.wobble 162.8 215.415 3.89778142024 4
pupil_05:34:51.wobble 161.6225 214.07 9.79921808105 4
pupil_04:15:22.wobble 162.07 214.2425 3.06220100581 4

  • SX M3 Pupil Wobble Plot:
    Screenshot at 2020-09-03 22-20-17.png

Collimation and Pointing Parameter update (after moving pupils around)

06:45 ACQUIRE preset to get new pointing on DX after the M3 update, SX didn't move M3 so only normal pointing drift (temperature is cooling fast).

06:55 ACTIVE preset to BS9188

Field Aberration Data on both sides

/home/lbto/data/20180907/dx_field and sx_field

BS9188 on-axis 0.8 hours east of meridian at EL=78. Seeing is fluctuating about 0.9 arcsec FWHM.

SX active optics ran away on first preset, but seemed to converge second time. We are collimating with GCS, but adjustment seems off a bit. Hopefully IDL can collimate well enough for the field aberration data.

DX active optics converges well, but pupil is one lenset too high on GCSGUI. This is because the pupil is slightly off the upper edge of the lenslet array in the on-axis position.

07:07 DX field_collect, limit=300 on BS9188 list - right WFSC display has off-axis pupil badly displaced, but GCS has it OK?

07:10 SX field_collect, limit=300 - collimation is struggling, maybe because of a wrinkle on the AdSec?

07:25 have lost star out of the acq field on both sides - too high elevation? Alas, pointing is still poor.

07:29 ACQUIRE preset for pointing update, but the pointing was still good 2 hours west of meridian. ?

07:38 readjusted SX pointing on BS9188 itself.

07:39 DX field_collect, limit=300 on BS9188 list

SX AGW4
                                                    IDL                                GCS
    Pupil Positioning:                  X         Y                       X           Y
          On-Axis Pupil =           170.70  218.60              170.70  218.60
                Pupil Shift =             0.00    0.01                    0.00    0.01
Shifted Pupil Center =           170.70  218.61
         Outer Pupil Fit =           160.75  213.22
                                              --------------                     --------------
            Pupil Center =           172.73  217.97              173.00  218.00
     
Shack-Hartmann Results:
         SH Spots Found =                115                          114
           SH Spots Used =                 67                           66
              Subaps Used =                 67                           68
Background Removed =                671                          671
X,Y Average Removed =           0.3342 -0.1161            0.2939 -0.1175
                 FWHM Ave =           0.70" +- 0.12"               0.68" +- 0.12"

07:46 SX field_collect, limit=300 on BS9188 list

DX lost the star - seems to be awful off-axis pointing, or a bad star in list - was gs=13

09:33 stopping collection of field data

20180907 Field Measurement Results DLM

 => Field Data Results
             Measured bi-nodal astigmatism
  
                  rfbg
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -25.2", Delta IE =  25.2"
         Tilt =   3.0", Delta CA =  -3.0"
  
  
                  rfbg
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -11.2", Delta IE =  11.2"
         Tilt =  -2.0", Delta CA =   2.0"
  
  
                  rfbg
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -15.2", Delta IE =  15.2"
         Tilt =  -3.2", Delta CA =   3.2"
  

NOTES DLM

The correlation of the 45 Astigmatism was better than the 0 Astigmatism and total astigmatism so I would suggest moving DX M2 by
  • Tip = -20", Delta IE = 20"
  • Tilt = 0", Delta CA = 0"

NOTES JMH Tip=-20 arcsec applied to M2 makes dX=0 dY=+1.021 dZ=0 dRX=-196.4 dRY=0
These were added permanently to DXSM Global Offsets on 20180912 UT - JMH

LBC

UT 09:34 switching to LBCs

UT 09:47 - Collimation and co-pointing near 2017 WV13 (Hazardous Asteroid!)

UT 09:56 - Moving to CP_2017_WV13_09062018.ob field for collimation

UT 10:03 - Running DOHYBRID

UT 10:14 - Running co-pointing

UT 10:18 - sending 2017WV13.xml script - non-sidereal guiding turned off in LBC GUI

UT 10:33 - 2017_WV13 finished


UT 10:34 - moving to specphot fields - Galadi 1

UT 10:35 - running DOFPIA

UT 10:41 - copointing completed

UT 10:42 - sending script for Galadi 1 all filters (airmass 1.108)

UT 11:02 - Galadi 1 done, sending SA92 co-point for collimation

UT 11:03 - Running DOFPIA

UT 11:08 - Running co-opointing

UT 11:12 - Running SA92 all filters (airmass 1.367)

UT 11:32 - SA92 done

UT 11:32 - moving to NGC 890 CP for collimation

UT 11:34 - running DOFPIA

UT 11:38 - Starting NGC 890 script (skipping co-pointing) - airmass 1.026

UT 12:07 - Stopping NGC 890 script at 4/5 dither position (4 Uspecs 8 I-Bessell images) - taking 1 short exposure (5sec) for Uspec and I-Bessell (for some reason it took 2x5sec each)


UT 12:10 - going to do Skyflats now

UT 12:15 - doing Uspec and I-Bessell (scale=8) PA=0

UT 12:22 - Uspec and I-Bessell (scale=2) PA=180

UT 12:27 - g-sloan and i-sloan (scale=1) PA=180 - 3/5 g-sloan OK, i-sloan probably all ok

UT12:32 - skyflats complete -sky too bright to continue

UT 12:33 - Closing Enclosure, powering off LBCs


-- %USERSIG{JohnHill - 2018-09-07}%
Topic attachments
I Attachment Action Size Date Who Comment
Screenshot at 2020-09-03 22-15-49.pngpng Screenshot at 2020-09-03 22-15-49.png manage 49 K 03 Sep 2020 - 22:26 JohnHill DX M3 Pupil Wobble Plot
Screenshot at 2020-09-03 22-20-17.pngpng Screenshot at 2020-09-03 22-20-17.png manage 44 K 03 Sep 2020 - 22:26 JohnHill SX M3 Pupil Wobble Plot
Topic revision: r9 - 03 Sep 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback